# SimSUM – Simulated Benchmark with Structured and Unstructured Medical Records

### v3

@@ Abstract
Methods: We introduce SimSUM [^1] An earlier version of this dataset was published under the name SynSUM. It has since been renamed to SimSUM to avoid confusion with synthetic data generated from real data, and to emphasize the simulated nature of the dataset., a benchmark dataset of 10,000 simulated patient records that link unstructured clinical notes with structured background variables. Each record simulates a patient encounter in the domain of respiratory diseases and includes tabular data (e.g., symptoms, diagnoses, underlying conditions) generated from a Bayesian network whose structure and parameters are defined by domain experts. A large language model (GPT-4o) is prompted to generate a clinical note describing the encounter, including symptoms and relevant context. These notes are annotated with span-level symptom mentions. We conduct an expert evaluation to assess note quality and run baseline predictive models on both the tabular and textual data.
@@ 1 Background
Figure 1: We have a clinical description of a patient encounter from which we want to extract some concepts, in this case the symptoms experienced by the patient. Some symptoms might be easy to extract using text-matching, like “high fever”. Other symptoms are not mentioned verbatim and are therefore harder to extract, like dyspnea. In this case, additional information on the patient, present in encoded format in the tabular portion of the EHR, together with domain knowledge, may help.
@@
Example 1: We know that the patient has asthma. Domain knowledge tells us that the probability of experiencing dyspnea when one has asthma is 90%. This can increase the confidence of the information extraction module for the concept of dyspnea being mentioned in the text.
@@
Example 2: We know that the patient is experiencing high fever. Domain knowledge tells us that a high fever often co-occurs with dyspnea due to their common cause, which is pneumonia. Even if we do not know that the patient has pneumonia, the probability of dyspnea being mentioned in the text increases as a result of observing high fever. If we model the joint probability of dyspnea, fever and pneumonia using a Bayesian network, we can get the exact probability of $\mathcal{P}(\text{{dyspnea = yes}}\mid\text{{fever = high}})$ by summing over the presence of pneumonia.
@@
Electronic health records (EHRs) are a gold mine of information, containing a mix of structured tabular variables (medication, diagnosis codes, lab results…) and free unstructured text (detailed clinical notes from physicians, nurses…) [4]. These EHRs form a valuable basis for training clinical decision support systems, (partially) automating essential processes in the clinical world, such as diagnosis, writing treatment plans, and more [22, 20, 25, 17, 32]. While large language models can help leverage the potential of the unstructured text portion of the EHR [33, 18, 8, 16, 29, 14], these black box systems lack interpretability [23, 34, 30]. In high-risk clinical applications, it can be argued that one should prefer more robust and transparent systems built on simpler, feature-based models, like regression models, decision trees, or Bayesian networks [27, 28, 19]. However, such models cannot directly deal with unstructured text and require tabular features as an input. For this reason, automated clinical information extraction (CIE) [4, 31, 6] is an essential tool for building large structured datasets that can serve as training data for such systems.
@@
Clinical information extraction remains a challenging task due to the complex nature of clinical notes. These often leave out important contextual details which an automated system would need in order to correctly extract concepts from the text. Existing systems do not fully exploit the available medical domain knowledge to fill in this gap. We propose that CIE could benefit from leveraging two additional sources of information, apart from the unstructured text itself. On the one hand, a range of tabular features are already encoded in the EHR. These contain information related to a particular patient visit (e.g. partially encoded symptoms or diagnosis codes), as well as information on the medical history of the patient. On the other hand, we can connect this encoded background information with the concepts we are trying to extract from the text, using a Bayesian network (BN) that represents medical domain knowledge. A visual example that helps illustrate this idea is shown in Fig. 1.
@@
To investigate and implement this idea, we need a clinical dataset which (i) contains a mix of tabular data and unstructured text, where (ii) the tabular data and the concepts we aim to extract from the text can be linked through domain knowledge. While open-source datasets like MIMIC-III [10] and MIMIC-IV [9] contain this mix, they are not a perfect fit. First, the area of intensive care in which the data was collected is very extensive, making it hard to isolate a specific use-case for which the domain knowledge could be listed. Second, the portion of the dataset which is encoded into tabular features is often driven by billing needs, rather than completeness or accuracy, and does not contain any encoded symptoms, which are concepts that could be interesting to extract from the text for application in clinical decision support systems. Third, the link between the tabular features and the concepts mentioned in the text might be inconsistent due to system design or human errors [13]. Finally, the EHRs in MIMIC being time series adds another layer of complexity. Other existing datasets linking unstructured clinical text to structured features include BioDEX [3], a large set of papers describing adverse drug events, as well as TCGA-Reports [11], a set of cancer pathology reports, both accompanied by tabular patient descriptors and extracted biological features. However, in both cases, it is not trivial to devise a BN representing the relevant expert knowledge, partially because that knowledge is not fully understood yet.
@@
In this work, we simulate a sufficiently realistic dataset that addresses some of these shortcomings, enabling research on incorporating domain knowledge for improved CIE in the presence of tabular variables. Our dataset, called SimSUM (Simulated Structured and Unstructured Medical records) is a self-contained set of artificial EHRs in a primary care setting, fulfilling the following requirements:
  • It mixes structured tabular data and unstructured text.
  • By design, clinical concepts expressed in the text and encoded in the tabular portion of the dataset are connected through a Bayesian network representing domain knowledge. In this case, the domain is respiratory diseases, with their associated symptoms and underlying conditions.
@@
  • Each EHR is a static snapshot of a single patient encounter, eliminating the time aspect for simplicity.
  • The text contains additional context on some of the encoded tabular variables.
@@
It is important to emphasize that the SimSUM dataset is not intended for training clinical decision support systems or deploying predictive models in real-world healthcare. Instead, it provides a controlled research environment for developing and prototyping methods. Although designed to be domain-relevant, SimSUM is entirely synthetic: tabular data is generated via an expert-defined Bayesian network, and clinical notes are produced by a large language model. It does not reflect the full variability or complexity of real EHRs, and should not be used to evaluate clinical model performance or train production models. Its value lies in enabling research on integrating background information into clinical information extraction, within a simplified setting with known and controllable ground-truth structure.
The remainder of this work is structured as follows. Section 2 describes how we generated SimSUM by simulating 10,000 patient encounters by sampling from an expert-defined BN and prompting a large language model. Section 3 analyses the utility of our dataset, and discusses its intended use. Finally, Section 4 states our main conclusions and highlights SimSUM’s limitations.
@@ 2 Construction and content
Figure 2: Overview of the full data generating process for the SimSUM dataset. First, the tabular portion of the artificial patient record is sampled from a Bayesian network, where both the structure and the conditional probability distributions were defined by an expert. Afterwards, we construct a prompt describing the symptoms experienced by the patient, as well as their underlying health conditions (but no diagnoses). We ask the GPT-4o large language model to generate a clinical note describing this patient encounter. Finally, we ask to generate a more challenging compact version of the note, mimicking the complexity of real clinical notes by prompting the use of abbreviations and shortcuts. We generate $10{,}000$ of these artificial patient records in total.
@@ 2.1 Modeling structured tabular variables with a Bayesian network
Asthma underlying condition Chronic lung disease in which the airways narrow and swell. yes/no
Smoking underlying condition Whether the patient is a regular smoker of tobacco. yes/no
COPD underlying condition Chronic Obstructive Pulmonary Disease, where airflow from the lungs is obstructed. yes/no
Hay fever underlying condition Allergic rhinitis, irritation of the nose caused by an allergen (e.g. pollen). yes/no
Season non-clinical Season of the year. winter/summer
Pneumonia diagnosis Infection that inflames the air sacs in one or both lungs. yes/no
Common cold diagnosis Upper respiratory tract infection, irritation and swelling of the upper airways. yes/no
Dyspnea symptom Shortness of breath, the feeling of not getting enough air. yes/no
Cough symptom Any type of cough. yes/no
Pain symptom Pain related to the airways or chest area. yes/no
Fever symptom Elevation of body temperature. high/low/none
Nasal symptom Nasal symptoms, such as runny nose or sneezing. yes/no
Policy non-clinical Whether the clinician has higher or lower prior inclination to prescribe antibiotics. Can be influenced by many factors, such as local policy in their general practice, their own caution towards antibiotics or level of experience. high/low
Self-employed non-clinical Whether the patient is self-employed, rendering them less inclined to stay home from work for longer periods. yes/no
Antibiotics treatment Whether any type of antibiotics are prescribed to the patient. yes/no
# Days at home outcome How many days the patient ends up staying home as a result of their symptoms and treatment. discrete $(0-...)$
Causal structure We asked an expert to define a Directed Acyclic Graph (DAG) which (partially) models the domain of respiratory diseases in primary care, shown in Fig. 2. In this DAG, a directed arrow between two variables models a causal relation between them. Central to the model are the diagnoses of pneumonia and common cold, which may give rise to five symptoms (dyspnea, cough, pain, fever and nasal). The expert also modeled some relevant underlying conditions which may render a patient more predisposed to certain diagnoses or symptoms: asthma, smoking, COPD and hay fever. Based on the symptoms experienced by a patient, a primary care doctor decides whether to prescribe antibiotics or not. The presence and severity of the symptoms, as well as the prescription of antibiotics as a treatment, influence the outcome, which is the total number of days (# days) that the patient eventually stays home as a result of illness. Finally, there are some non-clinical variables which exert an external influence on the diagnoses, treatment and outcome (season, policy and self-employed). Table 1 summarizes all variables and their meaning, as well as their possible values. While this model does not completely describe the real world, we do believe it to be sufficiently realistic for the purpose of generating a useful artificial dataset.
@@
$&\mathcal{P}_{\text{{joint}}}(\text{{asthma, smoking, ..., antibio, \# days}})=\\ &\mathcal{P}(\text{{asthma}})\mathcal{P}(\text{{smoking}})\mathcal{P}(\text{{COPD}}\mid\text{{smoking}})\mathcal{P}(\text{{hayf}})\mathcal{P}(\text{{season}})\mathcal{P}(\text{{cold}}\mid\text{{season}})\\ &\mathcal{P}(\text{{pneu}}\mid\text{{asthma, COPD, season}})\mathcal{P}(\text{{dysp}}\mid\text{{asthma, smoking, COPD, pneu, hayf}})\\ &\mathcal{P}(\text{{cough}}\mid\text{{asthma, smoking, COPD, pneu, cold}})\mathcal{P}(\text{{pain}}\mid\text{{cough, pneu, COPD, cold}})\\ &\mathcal{P}(\text{{fever}}\mid\text{{pneu, cold}})\mathcal{P}(\text{{nasal}}\mid\text{{cold, hayf}})\mathcal{P}(\text{{policy}})\mathcal{P}(\text{{self{\text{-}}empl}})\\ &\mathcal{P}(\text{{antibio}}\mid\text{{policy, dysp, cough, pain, fever}})\\ &\mathcal{P}(\text{{\#days}}\mid\text{{antibio, dysp, cough, pain, fever, nasal, self{\text{-}}empl}})\\ \end{aligned}$
We use four different approaches to parameterize these conditional distributions.
1. Conditional probability table When the variable is discrete and has a limited number of parents, we ask the expert to define a conditional probability table (CPT). Each entry of the table contains the probability for a particular value of the variable, conditional on the combination of values of the parent variables. If the variable has no parents, we just define a prior probability. The probabilities in the tables were filled in based on the expert’s own experience, as well as demographics in their local general practice and Belgium. While we do not expect these probabilities to generalize to the global patient population as a whole, a realistic-looking distribution suffices for our use-case. We provide these tables for the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever, policy and self-employed in Fig. 3.
@@
2. Noisy-OR distribution  For categorical variables with many parents, it becomes infeasible to manually fill in the CPT in a clinically meaningful way, because of the large number of possible combinations of parent values. This is the case for the symptoms dyspnea, cough, pain and nasal in our BN. To circumvent this problem, we define a Noisy-OR distribution, which assumes an independent causal mechanism behind the activation of a symptom through any of its parents [12]. This is a reasonable assumption to make in the case of symptoms with multiple possible causes (parents in the BN): a symptom arises in a patient if any of its possible causes succeeds in activating the symptom through its own independent mechanism. As shown in Equation (2), the parameterization of the Noisy-OR distribution rests on choosing each parameter $p_{i}$, which is the probability that a possible cause $X_{i}$ activates symptom $Y$. As a special case, $p_{0}$, also known as the leak probability, is the probability that symptom $Y$ is activated as the result of another unmodeled cause (outside of all $X_{i}$’s). Note that $x_{i}$ is $1$ when the cause $X_{i}$ is present in the patient (“yes” in our BN), and 00 if not.
@@
[…]
@@
$\begin{split}&\mathcal{P}(\text{{nasal}}\mid\text{{hayf, cold}})=NoisyOR(p_{0}=0.1,p_{\text{{hayf}}}=0.85,p_{\text{{cold}}}=0.7)\end{split}$ (6)
3. Logistic regression  We model the prescription of antibiotics by mimicking the way the clinician’s suspicion of pneumonia (which needs to be treated with antibiotics) rises when a higher number of symptoms is present in the patient, with some symptoms weighing more than others. Once their level of suspicion reaches a certain threshold, they decide to prescribe treatment. As shown in Equation (7), this process can be modeled using a logistic regression model taking as an input the symptoms dyspnea, cough, pain and fever, as well as the variable policy. Here, $x_{po}$ (policy) can take on the values 1 (high) or 0 (low), $x_{d}$ (dyspnea), $x_{c}$ (cough) and $x_{pa}$ (pain) can take on the value 1 (yes) or 0 (no), and $x_{f}$ (fever) can be 2 (high), 1 (low) or 0 (none).
@@
We decided on the weights in this model with help of the expert. First, we set the bias term and the coefficient for policy based on the following constraint: if there’s no symptoms at all, and policy is low, then the probability of prescribing antibiotics (due to some other unmodeled cause) should be around 5% . If policy is high, it should be around 10%. All other coefficients were then chosen by the expert based on the relative importance of the symptoms when deciding to prescribe antibiotics. As a final sanity-check, we asked the expert to label a set of test cases with whether or not they would prescribe antibiotics, and compared these with the probability predicted by the model. Table 7 in Appendix A shows these results. We see that the predictions made by the model mostly correspond well with the clinician’s intuition, confirming that the proposed coefficients make sense.
@@
4. Poisson regression  In a similar fashion, we model the # days the patient ends up staying home due to their complaints. Since the outcome is discrete, with most patients staying home for a low number of days, we use a Poisson regression. Assuming a non-linear effect of prescribing antibiotics, we define two separate models: one where antibiotics were not prescribed (Equation (8)), and one where they were (Equation (9)). Both models take as an input the symptoms dysp, cough, pain, nasal and fever, as well as the variable self-employed, and predict a mean number of days $\lambda$, which parameterizes the Poisson distribution.
@@
$\begin{split}&\mathcal{P}(\text{{\# days}}\mid\text{{dysp}}=x_{d},\text{{cough}}=x_{c},\text{{pain}}=x_{pa},\text{{nasal}}=x_{n},\text{{fever}}=x_{f},\\ &\text{{self{\text{-}}empl}}=x_{se},\text{{antibio}}=\text{{no}})=Poisson(\lambda_{0})\\ &\lambda_{0}=exp(0.010+0.64\times x_{d}+0.35\times x_{c}+0.47\times x_{pa}+0.011\times x_{n}\\ &+0.81\times(x_{f}==1)+1.23\times(x_{f}==2)-0.5\times x_{se})\\ &\text{with }x_{d},x_{c},x_{pa},x_{n},x_{se}\in\{0,1\}\text{, and }x_{f}\in\{0,1,2\}\end{split}$ (8)
@@
$\begin{split}&\mathcal{P}(\text{{\# days}}\mid\text{{dysp}}=x_{d},\text{{cough}}=x_{c},\text{{pain}}=x_{pa},\text{{nasal}}=x_{n},\text{{fever}}=x_{f},\\ &\text{{selfempl}}=x_{se},\text{{antibio}}=\text{{yes}})=Poisson(\lambda_{1})\\ &\lambda_{1}=exp(0.16+0.51\times x_{d}+0.42\times x_{c}+0.26\times x_{pa}+0.0051\times x_{n}\\ &+0.24\times(x_{f}==1)+0.57\times(x_{f}==2)-0.5\times x_{se})\\ &\text{with }x_{d},x_{c},x_{pa},x_{n},x_{se}\in\{0,1\}\text{, and }x_{f}\in\{0,1,2\}\end{split}$ (9)

### v4

@@ Abstract
Methods: We introduce SimSUM, a benchmark dataset of 10,000 simulated patient records that link unstructured clinical notes with structured background variables. Each record simulates a patient encounter in the domain of respiratory diseases and includes tabular data (e.g., symptoms, diagnoses, underlying conditions) generated from a Bayesian network whose structure and parameters are defined by domain experts. A large language model (GPT-4o) is prompted to generate a clinical note describing the encounter, including symptoms and relevant context. These notes are annotated with span-level symptom mentions. We conduct an expert evaluation to assess note quality and run baseline predictive models on both the tabular and textual data.
@@ 1 Background
Figure 1: General overview of the structure of this paper. We first describe the construction of SimSUM, a simulated dataset combining structured tabular data and unstructured clinical text (Section 2). We then evaluate the clinical notes through both a clinical expert review (Section 3.1) and automated span-level symptom extraction (Sections 2.3 and 3.2). Finally, we present four baseline predictive models that take as input the tabular data, the textual data, or both (Section 3.3).
@@
Figure 2: We have a clinical description of a patient encounter from which we want to extract some concepts, in this case the symptoms experienced by the patient. Some symptoms might be easy to extract using text-matching, like “high fever”. Other symptoms are not mentioned verbatim and are therefore harder to extract, like dyspnea. In this case, additional information on the patient, present in encoded format in the tabular portion of the EHR, together with domain knowledge, may help. We illustrate this idea with two examples.
@@
Example 1: We know that the patient has asthma. Domain knowledge may tell us that the probability of experiencing dyspnea when one has asthma (Equation (1)) is 90%, thereby increasing the prior probability of encountering dyspnea in the text. By integrating this knowledge in the information extraction module, it can more accurately predict the posterior probability of encountering dyspnea in the text.
@@
Example 2: We know that the patient is experiencing high fever. Domain knowledge may tell us that a high fever often co-occurs with dyspnea due to their common cause, which is pneumonia. Even if we do not know that the patient has pneumonia, the probability of dyspnea being mentioned in the text increases as a result of observing high fever. By modeling the joint probability of dyspnea, fever and pneumonia using a Bayesian network, we can get the exact probability of $\mathcal{P}(\text{{dyspnea = yes}}\mid\text{{fever = high}})$ by summing over the possible presence and absence of pneumonia in a procedure called Bayesian inference (Equation (2), [koller2009probabilistic]).
@@
Electronic health records (EHRs) are a gold mine of information, containing a mix of structured tabular variables [^1] In the remainder of our work, we use the terms “structured tabular data” and “tabular features” interchangeably. (medication, diagnosis codes, lab results…) and free unstructured text (detailed clinical notes from physicians, nurses…) [extracting_information_EMR]. These EHRs form a valuable basis for training clinical decision support systems, (partially) automating essential processes in the clinical world, such as diagnosis, writing treatment plans, and more [CDS_infectious_diseases, clinical_text_classification, medBERT, BEHRT, multimodal_ICD]. While large language models can help leverage the potential of the unstructured text portion of the EHR [combining_structured_unstructured, multimodal_matters, clinicalBERT, clinicalT5, MedPalm, BioMistral], these black box systems lack interpretability [black_box, explainability_LLMs, opportunities_challenges_biomedicine]. In high-risk clinical applications, it can be argued that one should prefer more robust and transparent systems built on simpler, feature-based models, like regression models, decision trees, or Bayesian networks [rudin2019stop, CML_for_healthcare, explainable_trees]. However, such models cannot directly deal with unstructured text and require tabular features as an input. For this reason, automated clinical information extraction (CIE) [extracting_information_EMR, clinical_information_extraction] is an essential tool for building large structured datasets that can serve as training data for such systems.
@@
Clinical information extraction remains a challenging task due to the complex nature of clinical notes [medical_information_extraction]. These often leave out important contextual details which an automated system would need in order to correctly extract concepts from the text. State-of-the-art clinical information extraction systems are predominantly built on pre-trained language models. Earlier methods relied on models pre-trained for entity and relation extraction in clinical text [pretrained_biomedical], while more recent approaches draw on large language models that have absorbed substantial medical knowledge during training [LLMs_information_extraction]. In both paradigms, domain knowledge is acquired implicitly rather than being explicitly encoded. Moreover, current systems typically overlook complementary information available in structured patient records. We propose that CIE could benefit from leveraging two additional sources of information, apart from the unstructured text itself. On the one hand, a range of tabular features are already encoded in the EHR. These contain information related to a particular patient visit (e.g. partially encoded symptoms or diagnosis codes), as well as information on the medical history of the patient. On the other hand, we can connect this encoded background information with the concepts we are trying to extract from the text, using a Bayesian network (BN) that represents medical domain knowledge. While ontologies and controlled vocabularies are commonly used to capture standardized concepts and semantic relationships [medical_informed_ML, medical_KG], they lack the ability to model causal dependencies or uncertainty. An expert-defined Bayesian network (BN) complements these resources by modeling how clinical factors influence one another and with what probabilistic strength [BN_healthcare]. A visual example that helps illustrate this idea is shown in Fig. 2.
@@
To investigate and implement this idea, we need a clinical dataset which (i) contains a mix of tabular data and unstructured text, where (ii) the tabular data and the concepts we aim to extract from the text can be linked through domain knowledge. While open-source datasets like MIMIC-III [mimic3] and MIMIC-IV [mimic4] contain this mix, they are not a perfect fit. First, the area of intensive care in which the data was collected is very extensive, making it hard to isolate a specific use-case for which the domain knowledge could be listed. Second, the portion of the dataset which is encoded into tabular features is often driven by billing needs [mimic4], rather than completeness or accuracy, and does not contain any encoded symptoms, which are concepts that could be interesting to extract from the text for application in clinical decision support systems. Third, the link between the tabular features and the concepts mentioned in the text might be inconsistent due to system design or human errors [EHRcon, improving_ehr]. Finally, the EHRs in MIMIC are time series, and while temporal information is a standard aspect of EHRs, it adds significant modeling complexity, making simpler datasets preferable for prototyping novel methods. Other existing datasets linking unstructured clinical text to structured features include BioDEX [biodex], a large set of papers describing adverse drug events, as well as TCGA-Reports [tcga_reports], a set of cancer pathology reports, both accompanied by tabular patient descriptors and extracted biological features. However, in both cases, it is not trivial to devise a BN representing the relevant expert knowledge, partially because that knowledge is not fully understood yet.
@@
In this work, we simulate a sufficiently realistic dataset that addresses some of these shortcomings, enabling research on incorporating domain knowledge for improved CIE in the presence of tabular variables. Our dataset, called SimSUM [^2] An earlier version of this dataset was published under the name SynSUM [synsum_old]. It has since been renamed to SimSUM to avoid confusion with synthetic data generated from real data, and to emphasize the simulated nature of the dataset. (Simulated Structured and Unstructured Medical records) is a self-contained set of artificial EHRs in a primary care setting, fulfilling the following requirements:
@@
  • Each record in SimSUM is an extract representing a single patient encounter, eliminating the time aspect for simplicity.
  • Each record is made up of structured tabular data and unstructured text.
  • By design, clinical concepts expressed in the text and encoded in the tabular portion of SimSUM are connected through a Bayesian network representing domain knowledge. In this case, the domain is respiratory diseases, with their associated symptoms and underlying conditions.
  • The unstructured text in SimSUM contains additional context on some of the encoded tabular variables. For example, symptoms are stored as structured tabular data, while the text elaborates on the nature of these symptoms, including their severity, onset, and other details.
@@
It is important to emphasize that the SimSUM dataset is not intended for training clinical decision support systems or deploying predictive models in real-world healthcare. Instead, it provides a controlled research environment for developing and prototyping methods. Although designed to be domain-relevant, SimSUM is entirely synthetic: tabular data is generated via an expert-defined Bayesian network, and clinical notes are produced by a large language model. It simulates a specific use-case in a limited domain – respiratory diseases in primary care – without modeling any temporal dynamics. SimSUM therefore does not reflect the full variability or complexity of real EHRs, and should not be used to evaluate clinical model performance or train production models. Its value lies in enabling research on integrating background information into clinical information extraction, within a simplified setting with known and controllable ground-truth structure.
The remainder of this work is structured as shown in Figure 1. In short, Section 2 describes how we generated SimSUM by simulating 10,000 patient encounters by sampling from an expert-defined BN and prompting a large language model. Section 3 analyses the utility of our dataset and discusses its intended use. Finally, Section 4 states our main conclusions and highlights SimSUM’s limitations.
@@ 2 Construction and content
Figure 3: Overview of the full data generating process for the SimSUM dataset. First, the tabular portion of the artificial patient record is sampled from a Bayesian network, where both the structure and the conditional probability distributions were defined by an expert. Afterwards, we construct a prompt describing the symptoms experienced by the patient, as well as their underlying health conditions (but no diagnoses). We ask the GPT-4o large language model to generate a clinical note describing this patient encounter. Finally, we ask to generate a more challenging compact version of the note, mimicking the complexity of real clinical notes by prompting the use of abbreviations and shortcuts. We generate 10,00010{,}000 of these artificial patient records in total.
@@ 2.1 Modeling structured tabular variables with a Bayesian network
For this part of the process, we enlisted an expert with over 30 years of experience in primary care. The expert has a background in developing and applying Bayesian networks in clinical settings and received support from the authors for any technical questions that arose.
Causal structure We asked the expert to define a Directed Acyclic Graph (DAG) which (partially) models the domain of respiratory diseases in primary care, shown in Fig. 3. In this DAG, a directed arrow between two variables models a causal relation between them. Central to the model are the diagnoses of pneumonia and common cold, which may give rise to five symptoms (dyspnea, cough, pain, fever and nasal). The expert also modeled some relevant underlying conditions which may render a patient more predisposed to certain diagnoses or symptoms: asthma, smoking, COPD and hay fever. Based on the symptoms experienced by a patient, a primary care doctor decides whether to prescribe antibiotics or not. The presence and severity of the symptoms, as well as the prescription of antibiotics as a treatment, influence the outcome, which is the total number of days that the patient eventually stays home as a result of illness (days at home). Finally, there are some non-clinical variables which exert an external influence on the diagnoses, treatment and outcome (season, policy and self-employed). Table 1 summarizes all variables and their meaning, as well as their possible values. While this model is a simplification of a real-world clinical process, we believe it is sufficiently realistic to fulfill our purpose of generating a simulated dataset.
@@
Asthma underlying condition
Chronic lung disease in which the airways narrow and swell.
yes/no
Smoking underlying condition
Whether the patient is a regular smoker of tobacco.
yes/no
COPD underlying condition
Chronic Obstructive Pulmonary Disease, where airflow from the lungs
is obstructed.
yes/no
Hay fever underlying condition
Allergic rhinitis, irritation of the nose caused by an allergen (e.g. pollen).
yes/no
Season non-clinical
Season of the year.
winter/summer
Pneumonia diagnosis
Infection that inflames the air sacs in one or both lungs.
yes/no
Common cold diagnosis
Upper respiratory tract infection, irritation and swelling of the upper
airways.
yes/no
Dyspnea symptom
Shortness of breath, the feeling of not getting enough air.
yes/no
Cough symptom
Any type of cough.
yes/no
Pain symptom
Pain related to the airways or chest area.
yes/no
Fever symptom
Elevation of body temperature.
high/low/none
Nasal symptom
Nasal symptoms, such as runny nose or sneezing.
yes/no
Policy non-clinical
Whether the clinician has higher or lower prior inclination to prescribe
antibiotics. Can be influenced by many factors, such as local policy in
their general practice, their own caution towards antibiotics or level of
experience.
high/low
Self-employed non-clinical
Whether the patient is self-employed, rendering them less inclined to
stay home from work for longer periods.
yes/no
Antibiotics treatment
Whether any type of antibiotics are prescribed to the patient.
yes/no
Days at home outcome
How many days the patient ends up staying home as a result of their
symptoms and treatment.
discrete (0)(0-...)
@@
[…]
We use four different approaches to parameterize these conditional distributions. For this process, we again rely on our expert primary care practitioner. Given that SimSUM is intended as a simplified, controlled benchmark rather than a reflection of real EHR variability, we find input from a single domain expert sufficient to guide the design of its probability distribution.
1. Conditional probability table When the variable is discrete and has a limited number of parents, we ask the expert to define a conditional probability table (CPT). We provide these tables for the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever, policy and self-employed in Figure 4. The probabilities in the tables were filled in based on the expert’s own experience, as well as demographics in their local general practice and Belgium. While we do not expect these probabilities to generalize to the global patient population as a whole, an expert-informed distribution which contains realistic elements suffices for our use-case.
@@
2. Noisy-OR distribution  For categorical variables with many parents, it becomes infeasible to manually fill in the CPT in a clinically meaningful way, because of the large number of possible combinations of parent values. This is the case for the symptoms dyspnea, cough, pain and nasal in our BN. To circumvent this problem, we define a Noisy-OR distribution, which assumes an independent causal mechanism behind the activation of a symptom through any of its parents [koller2009probabilistic]. This is a reasonable assumption to make in the case of symptoms with multiple possible causes (parents in the BN): a symptom arises in a patient if any of its possible causes succeeds in activating the symptom through its own independent mechanism [noisy_or, koller2009probabilistic]. As shown in Equation (2), the parameterization of the Noisy-OR distribution rests on choosing each parameter $p_{i}$, which is the probability that a possible cause $X_{i}$ activates symptom $Y$. As a special case, $p_{0}$, also known as the leak probability, is the probability that symptom $Y$ is activated as the result of another unmodeled cause (outside of all $X_{i}$’s). Note that $x_{i}$ is $1$ when the cause $X_{i}$ is present in the patient (“yes” in our BN), and 0 if not.
@@
$\begin{split}&\mathcal{P}(\text{{dysp}}\mid\text{{asthma, smoking, COPD, hayf, pneu}})=\operatorname{Noisy-OR}(p_{0}=0.05,\\ &p_{\text{{asthma}}}=0.9,p_{\text{{smoking}}}=0.3,p_{\text{{COPD}}}=0.9,p_{\text{{hayf}}}=0.2,p_{\text{{pneu}}}=0.3)\end{split}$ (3)
$\begin{split}&\mathcal{P}(\text{{cough}}\mid\text{{asthma, smoking, COPD, pneu, cold}})=\operatorname{Noisy-OR}(p_{0}=0.07,\\ &p_{\text{{asthma}}}=0.3,p_{\text{{smoking}}}=0.6,p_{\text{{COPD}}}=0.4,p_{\text{{pneu}}}=0.85,p_{\text{{cold}}}=0.7)\end{split}$ (4)
@@
$\begin{split}&\mathcal{P}(\text{{nasal}}\mid\text{{hayf, cold}})=\operatorname{Noisy-OR}(p_{0}=0.1,p_{\text{{hayf}}}=0.85,p_{\text{{cold}}}=0.7)\end{split}$ (6)
3. Logistic regression  We model the prescription of antibiotics by mimicking the way the clinician’s suspicion of pneumonia (which needs to be treated with antibiotics) rises when a higher number of symptoms is present in the patient, with some symptoms weighing more than others. Once their level of suspicion reaches a certain threshold, they decide to prescribe treatment. As shown in Equation (7), this process can be approximated in a simplified way through a logistic regression model, taking as an input the symptoms dyspnea, cough, pain and fever, as well as the variable policy. Here, $x_{po}$ (policy) can take on the values 1 (high) or 0 (low), $x_{d}$ (dyspnea), $x_{c}$ (cough) and $x_{pa}$ (pain) can take on the value 1 (yes) or 0 (no), and $x_{f}$ (fever) can be 2 (high), 1 (low) or 0 (none).
@@
We decided on the weights in this model with help of the expert. First, we set the bias term and the coefficient for policy based on the following constraint: if there’s no symptoms at all, and policy is low, then the probability of prescribing antibiotics (due to some other unmodeled cause) should be around 5% . If policy is high, it should be around 10%. All other coefficients were then chosen by the expert based on the relative importance of the symptoms when deciding to prescribe antibiotics. As a final sanity-check, we asked the expert to label a set of test cases with whether or not they would prescribe antibiotics, and compared these with the probability predicted by the model. Table 6 in Appendix A shows these results. We see that the predictions made by the model mostly correspond well with the desired input-output behavior, indicating that the proposed coefficients make sense.
@@
4. Poisson regression  In a similar fashion, we model the number of days at home as a result of the patient’s complaints with a Poisson regression model. Assuming a non-linear effect of prescribing antibiotics, we define two separate models: one where antibiotics were not prescribed (Equation (8)), and one where they were (Equation (9)). Both models take as an input the symptoms dysp, cough, pain, nasal and fever, as well as the variable self-employed, and predict a mean number of days $\lambda$, which parameterizes the Poisson distribution.
@@
$\begin{split}&\mathcal{P}(\text{{days at home}}\mid\text{{dysp}}=x_{d},\text{{cough}}=x_{c},\text{{pain}}=x_{pa},\text{{nasal}}=x_{n},\text{{fever}}=x_{f},\\ &\text{{self{\text{-}}empl}}=x_{se},\text{{antibio}}=\text{{no}})=Poisson(\lambda_{0})\\ &\lambda_{0}=exp(0.010+0.64\times x_{d}+0.35\times x_{c}+0.47\times x_{pa}+0.011\times x_{n}\\ &+0.81\times(x_{f}==1)+1.23\times(x_{f}==2)-0.5\times x_{se})\\ &\text{with }x_{d},x_{c},x_{pa},x_{n},x_{se}\in\{0,1\}\text{, and }x_{f}\in\{0,1,2\}\end{split}$ (8)
@@
$\begin{split}&\mathcal{P}(\text{{days at home}}\mid\text{{dysp}}=x_{d},\text{{cough}}=x_{c},\text{{pain}}=x_{pa},\text{{nasal}}=x_{n},\text{{fever}}=x_{f},\\ &\text{{selfempl}}=x_{se},\text{{antibio}}=\text{{yes}})=Poisson(\lambda_{1})\\ &\lambda_{1}=exp(0.16+0.51\times x_{d}+0.42\times x_{c}+0.26\times x_{pa}+0.0051\times x_{n}\\ &+0.24\times(x_{f}==1)+0.57\times(x_{f}==2)-0.5\times x_{se})\\ &\text{with }x_{d},x_{c},x_{pa},x_{n},x_{se}\in\{0,1\}\text{, and }x_{f}\in\{0,1,2\}\end{split}$ (9)